OpenStudy (anonymous):

elaluate the function for x=-1,1,2 f(x)=4 x 7^x

OpenStudy (anonymous):

Simply apply -1, 1, and 2 into your function. Is it like this?\[4\times7^{x}\]

OpenStudy (anonymous):

I guess it is. x = -1, f(x) = 4(7)^-1 = 4/7 x = 1, f(x) = 4(7)^1 = 28 x = 2, f(x) = 4(7)^2 = 196

OpenStudy (anonymous):

If you wonder about anything, ask it in this group. :)